OpenStudy (anonymous):

y varies inversely as twice x. When x = 3, y = 8. Find y when x = 2. y = 4 y = 6 y = 9 y = 12

OpenStudy (diyadiya):

y varies inversely with x therefore y=k/x

OpenStudy (diyadiya):

y= k/x plug in the value of x & y x =3 & y=8 and can you find out x now?

OpenStudy (anonymous):

this is hyperbolic trajectory

OpenStudy (diyadiya):

@rfig.khalil Please don't post only the final answer ,It against the rules of OS Please Read the Code of Conduct http://openstudy.com/code-of-conduct

OpenStudy (diyadiya):

@TLC Did you understand till that part ?

OpenStudy (anonymous):

so sorry i have no idea about this rule i apologize

OpenStudy (diyadiya):

Okay :) No problem!

OpenStudy (anonymous):

yeah i go it thanks